OpenStudy (anonymous):

can some one help me solve : csc^2xsecx/sec^2x+csc^2x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

best idea is to first rewrite in terms of sine and cosine. then we will have a bunch of algebra to do. it is then easiest to do the algebra if we write a = cos(x) and b = sin(x) so as not to be confused. ready?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[\frac{csc^(x)sec(x)}{sec^2(x)+csc^2(x)}\]

OpenStudy (anonymous):

that first csc is squared

OpenStudy (anonymous):

ok rewriting in terms of sine and cosine we get \[\frac{\frac{1}{sin^2(x)cos(x)}}{\frac{1}{cos^(x)}+\frac{1}{sin^2(x)}}\]

OpenStudy (anonymous):

satellite i need help with the question i posted earlier please

OpenStudy (anonymous):

yeah that was a typo. now replace cosine by a and sine by b so we can do the algebra unimpeded.

OpenStudy (anonymous):

k give me a minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{\frac{1}{b^2a}}{\frac{1}{a^2}+\frac{1}{b^2}}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

denominator is \[\frac{a^2+b^2}{a^2b^2}\]

OpenStudy (anonymous):

i mean that is just the denominator. numerator is still \[\frac{1}{b^2a}\]

OpenStudy (anonymous):

now we return to trig for a second. a = cosine and b = sine and \[\sin^2(x)+\cos^2(x)=1\]

OpenStudy (anonymous):

so our denominator is \[\frac{a^2+b^2}{a^2b^2}=\frac{1}{a^2b^2}\] because \[a^2+b^2=1\]

OpenStudy (anonymous):

yup i gotcha

OpenStudy (anonymous):

so we have in total a compound fraction that looks like \[\frac{\frac{1}{b^2a}}{\frac{1}{a^2b^2}}\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

invert and multiply to get \[\frac{1}{b^2a}\times \frac{a^2b^2}{1}\]

OpenStudy (anonymous):

cancel to get \[a\]

OpenStudy (anonymous):

yup i get it thank you so much

OpenStudy (anonymous):

so answer is ... \[cos(x)\]!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

btw it is much easier to work with a and b when you write this out. otherwise the algebra is hard to manipulate. only one trig step in this whole thing was \[\sin^2(x)+\cos^2(x)=1\] the rest was algebra

OpenStudy (anonymous):

thanks